News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 5 April 2002 8.00am EST

During the last week the loads were interrupted on 2 April due to an execution of the Science Instrument safing SCS 107. The SCS was executed by the OBC as a result of the EPHIN hardware fault flag being set. Analysis of both real-time and dump data showed no problems with the hardware and a procedure was run on 3 April to re-initialize the OBC EPHIN process and resume science. A similar event occurred last December and a software patch to increase two glitch counters was identified to reduce further spurious trips of SCS 107. The patch is expected to be tested and brought for approval for uplink next week. Observations of NGC 6397, BR1202-0725 and HD133880 were impacted and will be rescheduled in an upcoming schedule.

Of note last week was a Chandra image release of the Coma cluster of galaxies. The image shows a concentration of relatively cool gas around two of the larger galaxies that is sustained in a delicate energy balance. For details see: http://chandra.harvard.edu/photo/2002/0150/index.html.

The observing schedule for the next week is shown below and includes an observation of GX 3+1 coordinated with RXTE.

------------------------------------------
MSH 11-62              ACIS-S       Apr 7
AX J1325.8-3920        ACIS-S       Apr 8
HEN2-104               ACIS-S
GX 349+2               ACIS-S/HETG
GX 3+1                 ACIS-S/HETG  Apr 9
GK Per                 ACIS-S/HETG
Radiation Belts
RX J170930.2-263927    ACIS-S       Apr 10
RCS1417+5305           ACIS-I
NGC 4911               ACIS-I       Apr 11
SU Uma                 ACIS-S/HETG  Apr 12
Radiation Belts
SU Uma                 ACIS-S/HETG
Cyg X-1                ACIS-S/HETG  Apr 13 
4C 15.55 Field         ACIS-I
SDSS095151+594556      ACIS-S       Apr 14
PSS0955+5940           ACIS-S
FTM0738+2750           ACIS-S
Crab Nebula Halo       ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports